OpenStudy (anonymous):

Identify which equation is a quadratic equation. y = 3x + 4 y + x2 = (x − 2)(x + 7) y = −4x(x − 4) + 2x y + 3 = (x + 12)(x2 + 2)

jhonyy9 (jhonyy9):

so i think 3rd but can you tell me why ?

OpenStudy (anonymous):

no, how do you think its 3?

OpenStudy (anonymous):

well the 3rd one

jhonyy9 (jhonyy9):

so if you make the calcule there will be x^2 + ... =0 what sign that is quadratic equation ok ?

OpenStudy (anonymous):

yess

jhonyy9 (jhonyy9):

good luck bye

OpenStudy (mertsj):

The first one is a line

OpenStudy (mertsj):

2. y + x2 = (x − 2)(x + 7) y+x^2=x^2+5x-14 y=5x-14 That is linear also.

OpenStudy (mertsj):

y = −4x(x − 4) + 2x y=-4x^2+16x+2x y=-4x^2+18x That is quadratic because the highest exponent is 2

OpenStudy (mertsj):

y + 3 = (x + 12)(x2 + 2) y+3=x^3+12x^2+2x+24 That one is cubic because the highest exponent is 3

OpenStudy (anonymous):

thank you guys